OpenStudy (anonymous):

1.Show an equation and a solution for the problem. Crane A can unload a ship in 10 hours and crane B can unload it in 14 hours. How long will it take the two cranes to unload the ship working together? Round to the nearest tenth. 2.Show an equation and a solution for the problem. It takes John 20 hours to paint a building. It takes Sam 15 hours to paint the same building. How long will it take for them to paint the building if they work together, with Sam starting one hour later than John? Round to the nearest tenth. 3.Show an equation and a solution for the problem. How much water must be added to 12 L (liters) of a 40% solution of alcohol to obtain a 30% solution? 4.Show an equation and a solution. How much pure antifreeze must be added to 12 L of a 40% solution to obtain a 60% solution. (Remember that pure antifreeze is 100% = 1.) 5.Write an equation and a solution for the problem. A health food store sells oatmeal for $3.60 per pound and bran flakes for $4.80 per pound. How many pounds of each should be used to get a mixture of 30 pounds that sells for $4.00 a pound? 6.Show an equation and a solution for the problem. Arnold needs a 25% solution of nitric acid. He has 20 milliliters (ml) of a 30% solution. How many ml of a 15% solution should he add to obtain the required 25% solution?

OpenStudy (anonymous):

HELP

hero (hero):

If one person or machine can do a job in time \(m\) and the other person can do a job in \(n\) time, then together they can do the job in \(t\) time: \(\dfrac{m \times n}{m + n} = t\)

hero (hero):

In this case \(m = 10\) \(n = 14\)

hero (hero):

So together the machines can do the job in \(\dfrac{(10)(14)}{10 + 14} = t\) hours

OpenStudy (anonymous):

right

hero (hero):

\(\dfrac{140}{24} = \dfrac{35}{6}\) \(= \dfrac{30 + 5}{6}\) \(=\dfrac{30}{6} + \dfrac{5}{6}\) \(=5 + \dfrac{5}{6}\) Hrs

hero (hero):

Or 5 hours and 50 minutes

hero (hero):

The second question is similar to the first question.

hero (hero):

So try that one on your own.

hero (hero):

By the way, I got 50 minutes by converting hours to minutes using this proportion: \(\dfrac{5}{6} = \dfrac{x}{60}\)

hero (hero):

The answer is correct, however, perhaps your program is expecting a less than exact answer since it says "round to the nearest tenth".